Satellite A orbits a planet at a distance d from the planet’s center with a centripetal acceleration a0. A second identical satellite B orbits the same planet at a distance 2d from the planet’s center with centripetal acceleration ab. What is the centripetal acceleration ab in terms of a0 ?
1
Expert's answer
2019-11-15T10:23:25-0500
"a_0=\frac{GM}{d^2}"

"a_b=\frac{GM}{(2d)^2}=\frac{GM}{4d^2}=\frac{a_0}{4}"